# Reconstrucion of oriented matroids from Varchenko-Gelfand algebras

### v1

@@ 4.1. Generalized Heaviside functions
$\widetilde{\mathcal{H}}(\mathcal{A})=\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))\setminus\{0,1\},$ (4.1)
and call an element $y\in\widetilde{\mathcal{H}}(\mathcal{A})$ a generalized Heaviside function. The key step in Section 3.2 (Lemma 3.8) was that if $\mathcal{A}$ is generic in codimension $2$, then
$\mathcal{H}(\mathcal{A})=\widetilde{\mathcal{H}}(\mathcal{A}).$ (4.2)
@@
$\operatorname{Alt}^{+}(H_{1},\dots H_{2s+1})$= $1_{C_{1}}+1_{C_{3}}+1_{C_{5}}+\cdots+1_{C_{4s+1}}$ (4.3)
@@ 4.2. Conjectural algorithm
By Theorem 3.6, 4.3 holds for arrangements that are generic in codimension $2$ when $\operatorname{char}R\neq 2$. As seen in Section 4.1, however, there are many generalized Heaviside functions that are not Heaviside functions. To formulate a reconstruction procedure from $\operatorname{\mathcal{VG}}(\mathcal{A})$, we introduce the notion of “generalized tope graphs.”
@@
The point of 4.5 is that even when $y_{1},\dots,y_{n}$ are not Heaviside functions, if the resulting graph is an oriented matroid tope graph, then it must be isomorphic to $\operatorname{\mathcal{T}}(\mathcal{A})$.
@@ 5.1. Square-zero elements
$\operatorname{\mathsf{N}^{2}}(\mathcal{A}):=\{\overline{y}\in\operatorname{\mathsf{VG}}^{1}(\mathcal{A})\mid\overline{y}^{2}=0\mbox{ in }\operatorname{\mathsf{VG}}^{2}(\mathcal{A})\}.$ (5.1)
@@
$\operatorname{\mathsf{N}^{2}}(\mathcal{A})=\bigcup_{y\in\widetilde{\mathcal{H}}(\mathcal{A})}R\cdot\overline{y}.$ (5.2)
@@
$\operatorname{\mathsf{N}^{2}}(\mathcal{A})=\{y\in\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})\mid y^{2}\in\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})\}/\operatorname{\mathsf{F}}^{0}.$ (5.3)
@@
$\alpha_{1}-\beta_{1}=\alpha_{2}-\beta_{2}=\cdots=\alpha_{s}-\beta_{s}=\gamma_{1}$ (5.4)
@@
$\alpha_{1}^{2}-\beta_{1}^{2}=\alpha_{2}^{2}-\beta_{2}^{2}=\cdots=\alpha_{s}^{2}-\beta_{s}^{2}=\gamma_{2}.$ (5.5)
@@
$\alpha_{1}+\beta_{1}=\alpha_{2}+\beta_{2}=\cdots=\alpha_{s}+\beta_{s}=\gamma_{3},$ (5.6)
@@
$\alpha_{1}=\alpha_{2}=\cdots=\alpha_{s}=\frac{\gamma_{1}+\gamma_{3}}{2},\ \ \beta_{1}=\beta_{2}=\cdots=\beta_{s}=\frac{\gamma_{3}-\gamma_{1}}{2}.$ (5.7)
@@
$\operatorname{\mathsf{N}^{2}}(\mathcal{A}_{1})=$ $\bigcup_{i=1}^{6}Rx_{i}\cup R(x_{1}+x_{2}-x_{3})\cup R(x_{1}-x_{4}+x_{5})$ (5.8)
@@ 5.2. Recovering signed circuits
$\operatorname{\mathsf{N}^{2}}(\mathcal{A})=R\cdot f_{1}\cup\cdots\cup R\cdot f_{n},$ (5.9)
@@
$R\cdot f_{i}=R\cdot\overline{x_{i}^{+}},$ (5.10)
@@
$\overline{x_{i_{1}}^{+}}\cdot\overline{x_{i_{2}}^{+}}\cdots\overline{x_{i_{k}}^{+}}\neq 0,\mbox{ in }\ \operatorname{\mathsf{VG}}^{k}(\mathcal{A}),$ (5.11)
@@
$f_{I}=0,\mbox{ and }f_{I\setminus\{i_{p}\}}\neq 0\mbox{ for $p=1,\dots,k$},$ (5.12)
@@
$\lambda^{I}_{1}\cdot f_{I\setminus\{i_{1}\}}+\lambda^{I}_{2}\cdot f_{I\setminus\{i_{2}\}}+\cdots+\lambda^{I}_{k}\cdot f_{I\setminus\{i_{k}\}}=0,$ (5.13)
@@
$\lambda_{i}=\begin{cases}\sigma_{p},\ \mbox{$i\in I$ and $i=i_{p}$}\\ 0,\ i\notin I.\end{cases}$ (5.14)
@@
$\check{\operatorname{\mathcal{C}}}(\overline{x_{1}^{+}},\dots,\overline{x_{n}^{+}})=\operatorname{\mathcal{C}}(\alpha_{1},\dots,\alpha_{n}).$ (5.16)
@@
$\lambda_{1}^{I}\cdot f_{I\setminus\{i_{1}\}}+\lambda_{2}^{I}\cdot f_{I\setminus\{i_{2}\}}+\cdots+\lambda_{k}^{I}\cdot f_{I\setminus\{i_{k}\}}=0$ (5.17)
$\lambda_{1}^{{}^{\prime}I}\cdot f_{I\setminus\{i_{1}\}}^{\prime}+\lambda_{2}^{{}^{\prime}I}\cdot f_{I\setminus\{i_{2}\}}^{\prime}+\cdots+\lambda_{k}^{{}^{\prime}I}\cdot f_{I\setminus\{i_{k}\}}^{\prime}=0,$ (5.18)
@@
$\mu_{I\setminus\{i_{1}\}}\lambda_{1}^{I}\cdot f^{\prime}_{I\setminus\{i_{1}\}}+\mu_{I\setminus\{i_{2}\}}\lambda_{2}^{I}\cdot f^{\prime}_{I\setminus\{i_{2}\}}+\cdots+\mu_{I\setminus\{i_{k}\}}\lambda_{k}^{I}\cdot f^{\prime}_{I\setminus\{i_{k}\}}=0.$ (5.19)
@@
$(\mu_{I\setminus\{i_{1}\}}\lambda_{1}^{I},\mu_{I\setminus\{i_{2}\}}\lambda_{2}^{I},\dots,\mu_{I\setminus\{i_{k}\}}\lambda_{k}^{I})=c_{I}\cdot(\lambda_{1}^{{}^{\prime}I},\lambda_{2}^{{}^{\prime}I},\dots,\lambda_{k}^{{}^{\prime}I}).$ (5.20)
@@
$\mu_{I\setminus\{i_{q}\}}=\frac{\lambda_{p}^{I}\cdot\lambda_{q}^{{}^{\prime}I}}{\lambda_{p}^{{}^{\prime}I}\cdot\lambda_{q}^{I}}\cdot\mu_{I\setminus\{i_{p}\}}=\pm\mu_{I\setminus\{i_{p}\}}.$ (5.21)
@@
$i\sim j\Longleftrightarrow\mbox{there exists a circuit $I\subset[n]$ such that $i,j\in I$}.$ (5.22)
@@
$\lambda_{1}^{I}\cdot f_{I\setminus\{i_{1}\}}+$ $\cdots+\lambda_{k}^{I}\cdot f_{I\setminus\{i_{k}\}}$ (5.24)
@@
Let $\mathcal{A}=\{H_{1},\dots,H_{n}\}$ be a real arrangement in $\mathbb{R}^{\ell}$. In analogy with Section 4.2, we propose a conjectural algorithm recovering $\operatorname{\mathcal{C}}(\mathcal{A})$ from $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})$ as follows.
@@
We conjecture that the resulting oriented matroid is equivalent to the original one. This procedure terminates in finitely many steps whenever there are only finitely many choices in (Step 1), e.g., the case $R^{\times}<\infty$. Unlike the filtered case (4.5), there can be infinitely many possibilities in (Step 1) when $R^{\times}=\infty$, so termination is not guaranteed in full generality.
@@ 5.3. Concluding remarks
In 4.3, we assumed $\operatorname{char}R\neq 2$. Here, we briefly discuss the case $\operatorname{char}R=2$. If $\operatorname{char}R=2$, we can not recover the oriented matroid from the graded VG algebra. The counterexample can be obtained again by six generic planes. Let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be six generic planes in $\mathbb{R}^{3}$ as in Example 3.10. Then $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})\simeq\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})$ when $\operatorname{char}R=2$, because
@@ Acknowledgements
The second author is supported by JSPS KAKENHI JP23H00081.

### v2

@@ 4.1. Generalized Heaviside functions
(4.1) $\widetilde{\mathcal{H}}(\mathcal{A})=\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))\setminus\{0,1\},$
and call an element $y\in\widetilde{\mathcal{H}}(\mathcal{A})$ a generalized Heaviside function. The key step in §3.2 (Lemma 3.8) was that if $\mathcal{A}$ is generic in codimension $2$, then
(4.2) $\mathcal{H}(\mathcal{A})=\widetilde{\mathcal{H}}(\mathcal{A}).$
@@
(4.3) $\operatorname{Alt}^{+}(H_{1},\dots H_{2s+1})$= $1_{C_{1}}+1_{C_{3}}+1_{C_{5}}+\cdots+1_{C_{4s+1}}$
@@ 4.2. Conjectural algorithm
By Theorem 3.6, Conjecture 4.3 holds for arrangements that are generic in codimension $2$ when $\operatorname{char}R\neq 2$. As seen in §4.1, however, there are many generalized Heaviside functions that are not Heaviside functions. To formulate a reconstruction procedure from $\operatorname{\mathcal{VG}}(\mathcal{A})$, we introduce the notion of “generalized tope graphs.”
@@
The point of Conjecture 4.5 is that even when $y_{1},\dots,y_{n}$ are not Heaviside functions, if the resulting graph is an oriented matroid tope graph, then it must be isomorphic to $\operatorname{\mathcal{T}}(\mathcal{A})$.
@@ 5.1. Square-zero elements
(5.1) $\operatorname{\mathsf{N}^{2}}(\mathcal{A}):=\{\overline{y}\in\operatorname{\mathsf{VG}}^{1}(\mathcal{A})\mid\overline{y}^{2}=0\mbox{ in }\operatorname{\mathsf{VG}}^{2}(\mathcal{A})\}.$
@@
(5.2) $\operatorname{\mathsf{N}^{2}}(\mathcal{A})=\bigcup_{y\in\widetilde{\mathcal{H}}(\mathcal{A})}R\cdot\overline{y}.$
@@
(5.3) $\operatorname{\mathsf{N}^{2}}(\mathcal{A})=\{y\in\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})\mid y^{2}\in\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})\}/\operatorname{\mathsf{F}}^{0}.$
@@
(5.4) $\alpha_{1}-\beta_{1}=\alpha_{2}-\beta_{2}=\cdots=\alpha_{s}-\beta_{s}=\gamma_{1}$
@@
(5.5) $\alpha_{1}^{2}-\beta_{1}^{2}=\alpha_{2}^{2}-\beta_{2}^{2}=\cdots=\alpha_{s}^{2}-\beta_{s}^{2}=\gamma_{2}.$
@@
(5.6) $\alpha_{1}+\beta_{1}=\alpha_{2}+\beta_{2}=\cdots=\alpha_{s}+\beta_{s}=\gamma_{3},$
@@
(5.7) $\alpha_{1}=\alpha_{2}=\cdots=\alpha_{s}=\frac{\gamma_{1}+\gamma_{3}}{2},\ \ \beta_{1}=\beta_{2}=\cdots=\beta_{s}=\frac{\gamma_{3}-\gamma_{1}}{2}.$
@@
(5.8) $\operatorname{\mathsf{N}^{2}}(\mathcal{A}_{1})=$ $\bigcup_{i=1}^{6}Rx_{i}\cup R(x_{1}+x_{2}-x_{3})\cup R(x_{1}-x_{4}+x_{5})$
@@ 5.2. Recovering signed circuits
(5.9) $\operatorname{\mathsf{N}^{2}}(\mathcal{A})=R\cdot f_{1}\cup\cdots\cup R\cdot f_{n},$
@@
(5.10) $R\cdot f_{i}=R\cdot\overline{x_{i}^{+}},$
@@
(5.11) $\overline{x_{i_{1}}^{+}}\cdot\overline{x_{i_{2}}^{+}}\cdots\overline{x_{i_{k}}^{+}}\neq 0,\mbox{ in }\ \operatorname{\mathsf{VG}}^{k}(\mathcal{A}),$
@@
(5.12) $f_{I}=0,\mbox{ and }f_{I\setminus\{i_{p}\}}\neq 0\mbox{ for $p=1,\dots,k$},$
@@
(5.13) $\lambda^{I}_{1}\cdot f_{I\setminus\{i_{1}\}}+\lambda^{I}_{2}\cdot f_{I\setminus\{i_{2}\}}+\cdots+\lambda^{I}_{k}\cdot f_{I\setminus\{i_{k}\}}=0,$
@@
(5.14) $\lambda_{i}=\begin{cases}\sigma_{p},\ \mbox{$i\in I$ and $i=i_{p}$}\\ 0,\ i\notin I.\end{cases}$
@@
(5.16) $\check{\operatorname{\mathcal{C}}}(\overline{x_{1}^{+}},\dots,\overline{x_{n}^{+}})=\operatorname{\mathcal{C}}(\alpha_{1},\dots,\alpha_{n}).$
@@
(5.17) $\lambda_{1}^{I}\cdot f_{I\setminus\{i_{1}\}}+\lambda_{2}^{I}\cdot f_{I\setminus\{i_{2}\}}+\cdots+\lambda_{k}^{I}\cdot f_{I\setminus\{i_{k}\}}=0$
(5.18) $\lambda_{1}^{{}^{\prime}I}\cdot f_{I\setminus\{i_{1}\}}^{\prime}+\lambda_{2}^{{}^{\prime}I}\cdot f_{I\setminus\{i_{2}\}}^{\prime}+\cdots+\lambda_{k}^{{}^{\prime}I}\cdot f_{I\setminus\{i_{k}\}}^{\prime}=0,$
@@
(5.19) $\mu_{I\setminus\{i_{1}\}}\lambda_{1}^{I}\cdot f^{\prime}_{I\setminus\{i_{1}\}}+\mu_{I\setminus\{i_{2}\}}\lambda_{2}^{I}\cdot f^{\prime}_{I\setminus\{i_{2}\}}+\cdots+\mu_{I\setminus\{i_{k}\}}\lambda_{k}^{I}\cdot f^{\prime}_{I\setminus\{i_{k}\}}=0.$
@@
(5.20) $(\mu_{I\setminus\{i_{1}\}}\lambda_{1}^{I},\mu_{I\setminus\{i_{2}\}}\lambda_{2}^{I},\dots,\mu_{I\setminus\{i_{k}\}}\lambda_{k}^{I})=c_{I}\cdot(\lambda_{1}^{{}^{\prime}I},\lambda_{2}^{{}^{\prime}I},\dots,\lambda_{k}^{{}^{\prime}I}).$
@@
(5.21) $\mu_{I\setminus\{i_{q}\}}=\frac{\lambda_{p}^{I}\cdot\lambda_{q}^{{}^{\prime}I}}{\lambda_{p}^{{}^{\prime}I}\cdot\lambda_{q}^{I}}\cdot\mu_{I\setminus\{i_{p}\}}=\pm\mu_{I\setminus\{i_{p}\}}.$
@@
(5.22) $i\sim j\Longleftrightarrow\mbox{there exists a circuit $I\subset[n]$ such that $i,j\in I$}.$
@@
(5.24) $\lambda_{1}^{I}\cdot f_{I\setminus\{i_{1}\}}+$ $\cdots+\lambda_{k}^{I}\cdot f_{I\setminus\{i_{k}\}}$
@@
Let $\mathcal{A}=\{H_{1},\dots,H_{n}\}$ be a real arrangement in $\mathbb{R}^{\ell}$. In analogy with §4.2, we propose a conjectural algorithm recovering $\operatorname{\mathcal{C}}(\mathcal{A})$ from $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})$ as follows.
@@
We conjecture that the resulting oriented matroid is equivalent to the original one. This procedure terminates in finitely many steps whenever there are only finitely many choices in (Step 1), e.g., the case $R^{\times}<\infty$. Unlike the filtered case (Conjecture 4.5), there can be infinitely many possibilities in (Step 1) when $R^{\times}=\infty$, so termination is not guaranteed in full generality.
@@ 5.3. Concluding remarks
In Conjecture 4.3, we assumed $\operatorname{char}R\neq 2$. Here, we briefly discuss the case $\operatorname{char}R=2$. If $\operatorname{char}R=2$, we can not recover the oriented matroid from the graded VG algebra. The counterexample can be obtained again by six generic planes. Let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be six generic planes in $\mathbb{R}^{3}$ as in Example 3.10. Then $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})\simeq\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})$ when $\operatorname{char}R=2$, because
@@ Acknowledgements
The authors thank Professor Takuro Abe for pointing out several mistakes in the previous version. The second author is supported by JSPS KAKENHI JP23H00081.